OpenStudy (anonymous):

This integral is giving me trouble (x^2 +2)/(x^2 +1) dx. Anyone want to be a hero and help me out?

OpenStudy (anonymous):

i think you have to divide first, then use partial fractions maybe i am wrong

OpenStudy (anonymous):

= ( x² + 1 ) / ( x² + 1 ) + 1/ ( x² + 1 ) Then take integral

OpenStudy (anonymous):

oh i am wrong but you do need to divide first, and there is an easy trick to doing it \[\frac{x^2+2}{x^2+1}=\frac{x^2+1+1}{x^2+1}=1+\frac{1}{x^2+1}\]

OpenStudy (anonymous):

what @Chlorophyll said

OpenStudy (anonymous):

btw you are supposed to recognize the second integral, not compute anything

OpenStudy (anonymous):

Wow, thank you so much guys. I was doing all kinds of crazy trig triangle subs, integration by parts, and u subs trying to figure this out. Glad to see it's so simple. Cheers!